OpenStudy (anonymous):

Chance can wash the car in 2 hours, while Sandy can wash the same car in 4 hours. How many hours would it take for them to wash the car together?

OpenStudy (anonymous):

is it 3/4 hours?

OpenStudy (ddcamp):

\[\frac{ 1 }{ Time\ for\ Chance } + \frac{ 1 }{ Time\ for\ Sandy } = \frac{ 1 }{ Time\ together }\]

OpenStudy (anonymous):

okay i think i got it